If a 120-W lightbulb emits 2.5 % of the input energy as visible light (average wavelength 550 nm) uniformly in all directions.

Part A

How many photons per second of visible light will strike the pupil (4.0 mm diameter) of the eye of an observer 1.6 m away?

Express your answer using two significant figures.


Part B

How many photons per second of visible light will strike the pupil (4.0 mm diameter) of the eye of an observer 1.5 km away?

Express your answer using two significant figures.If a 120-W lightbulb emits 2.5 % of the input energy as visible light (average wavelength 550 nm) uniformly in all directions
Answer #1

Solution:

Given :

P = 120 W ; % = 2.5 ;

Wavelength (\lambda) = 550 nm

Part (A) Solution:

Here, d = 4 mm ; D = 1.6 m

P = 0.025 x 120 = 3 W

We know that the energy carried by a single photon is:

E = h c / \lambda = (6.626 x 10-34)(3 x 108) / (550 x 10-9) = 3.61 x 10-19 J

We know that, intensity is:

I = P/A = 3 / { 4 x 3.14 x (1.6)2 } = 0.09325 W/m2

Therefore: at eye ; P = I x A

P = (0.09325 W/m2)(3.14) x (2 x 10-3)2 = 1.172 x 10-6 J/s

Thus: N = ( 1.172 x 10-6 J/s) / (3.61 x 10-19) = 3.246 x 1012 photons/s

Hence, N = 3.246 x 1012 photons/s

Part (B) Solution:

Here, I = P/A = (3 W) / {4 x 3.14 x (1500 m)2} = 1.1 x 10-7 W/m2

P = I x A (at eye)

P = 1.1 x 10-7 x 3.14 x (2 x 10-3)2 = 1.4 x 10-12 J/s

N = (1.4 x 10-12 J/s) /(3.61 x 10-19) = 3.9 x 106 photons/s

Hence, N = 3.9 x 106 photons/s